OpenStudy (anonymous):

MEDAL FOR RIGHT ANSWER!!! In a candy bar, 165 calories are from sugar. There are 250 calories in the candy bar. What percent of the calories are from sugar?

OpenStudy (anonymous):

@horsegirl27 @mathmate

OpenStudy (anonymous):

so basically, what percent of 250 is 165

OpenStudy (displayerror):

Let x represent the percent of calories that is from sugar. We can write the expression as \[x \times 250 = 165\] As you mentioned, this is the same as "what percent of (keep in mind "of" suggests multiplication) 250 is equal to 165?" Solve for x and multiply by 100 and you get your percent. \[\frac{x \times 250}{250} = \frac{165}{250}\] \[x = \frac{165}{250}\]

OpenStudy (horsegirl27):

^^^go with that